In triangle ABC, the measure of angle B is 5° more than three times the measure of angle A. The measure of angle C is 15° more than the measure of angle A. Which one of the following is true?
A) Angle A is 35° B) Angle B is 102° C) Angle C is 47° D) None of these

Answer by Boreal(15235)

A is x
B is 3x+5
C=x+15
they add to 180 degrees
so 5x+20=180
x=32 degrees A
3x+5=101 degrees B
x+15=47 degrees C
C